OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line? Answer -3/5 1/3 3 -3

OpenStudy (whpalmer4):

Solve your equation for \(y\) to get it in slope-intercept form: \(y = mx+b\) where \(m\) is slope. Now divide -1 by \(m\) to get the slope of the perpendicular line. The product of the slopes of perpendicular lines should equal -1 (unless one of them is vertical)

OpenStudy (anonymous):

How would you divide m?

OpenStudy (whpalmer4):

\(m\) is going to be a number. I'll walk you through it. Solve the equation for \(y\): what do you get?

OpenStudy (anonymous):

What is y again is the answer -1

OpenStudy (whpalmer4):

Come on, you have the equation \[x +3y = 5\] Solve that for \(y\). In other words, arrange it so \(y = \) <something>

OpenStudy (whpalmer4):

Do you know how to do that?

OpenStudy (whpalmer4):

\[x+3y=5\]Subtract \(x\) from both sides: \[x-x+3y = 5-x\]\[3y=5-x\]Divide both sides by 3: \[\frac{3y}{3} = \frac{5-x}{3}\]\[y=-\frac{1}{3}x+\frac{5}{3}\]

OpenStudy (whpalmer4):

Now, compare that with slope intercept form: \[y=mx+b\] Matching up the respective pieces, what is the value of \(m\)?

OpenStudy (anonymous):

ok would you add 1/3x + 5/3 to get m?

OpenStudy (whpalmer4):

\[y = mx + b\] \[y = -\frac{1}{3}x + \frac{5}{3}\] Match up the pieces! \(y\) matches \(y\). what matches with \(mx\)?

OpenStudy (anonymous):

-1/3

OpenStudy (whpalmer4):

Yes, \[m = -\frac{1}{3}\] Okay, the slope of the other line is \[\frac{1}{m} = \frac{1}{-\frac{1}{3}}=\] Do you know how to divide fractions?

OpenStudy (anonymous):

do you flip the fraction and multiply it?

OpenStudy (whpalmer4):

Flip the fraction in the denominator and multiply by it, yes.

OpenStudy (anonymous):

so that would be 1/m times -1/3/1

OpenStudy (anonymous):

-1/3/m

OpenStudy (whpalmer4):

no. we are trying to divide 1 by -1/3. 1 = 1/1 invert -1/3 -> -3/1 1/1 * -3/1 = -3

OpenStudy (whpalmer4):

Ah, @#$#$, we wanted to divide -1 by m, not 1. Sigh. \[\frac{-1}{m} = \frac{-1}{-\frac{1}{3}}\]to do that, we multiply the numerator (-1) by the inverted denominator (-3/1) \[-1*\frac{-3}{1} = 3\]

OpenStudy (anonymous):

okay that makes sense I did not know that

OpenStudy (whpalmer4):

We need the two slopes to multiply together and give -1 as the answer. Let's see if they do: \[-\frac{1}{3}*3=-\frac{3}{3} = -1\]They do, so 3 is the correct answer for the slope of the perpendicular line.

OpenStudy (whpalmer4):

Sorry for the confusion due to the missing - sign! This is an example of why I always check my work — I catch silly mistakes like this!

OpenStudy (anonymous):

okay great tip " to always check your work" Thank you so much for the help

OpenStudy (whpalmer4):

Hope it helped!

OpenStudy (jhannybean):

Another thing - slope of a perpendicular line follows the format \[m= -\frac{1}{m}\] where -1/m is the slope of the perpendicular line and m is the slope of the line given/found.

OpenStudy (dan815):



OpenStudy (dan815):